OpenStudy (hijacktrolo):

What is the value of y in the equation 2(2y − 16) = 0? 4 6 8 9 i think its 8 but want to be sure

OpenStudy (pinklion23):

yes you are correct :)

OpenStudy (hijacktrolo):

:D

OpenStudy (hijacktrolo):

thx

OpenStudy (hijacktrolo):

one more clarifying question

OpenStudy (hijacktrolo):

The work of a student to solve the equation 4(2x − 4) = 8 + 2x + 8 is shown below: Step 1: 4(2x − 4) = 8 + 2x + 8 Step 2: 6x − 8 = 16 + 2x Step 3: 6x − 2x = 16 + 8 Step 4: 4x = 24 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 8x − 4 = 2(6 + x + 6) Step 2; 8x − 16 = 16 + 2x Step 3; 6x − 2x = 16 − 8 Step 3; 6x + 2x = 16 + 8

OpenStudy (hijacktrolo):

i think its c

OpenStudy (pinklion23):

i think you are right

OpenStudy (hijacktrolo):

:D